# Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v2

@@ 2.4 Specifically Trained LLMs
The transformer model and the self-attention mechanism have paved the way for researchers to train language models on trillions of tokens with billions of parameters. Several LLMs have been trained and harnessed, each tailored with specific capabilities for diverse fields such as security, chemistry, engineering, medicine, business, tourism, and language-related applications. These models are employed in tasks ranging from detecting security threats, analysing data, and generating synthetic actions to teaching, code generation, structured query generation, planning, assisting in medical education, clinical decision-making, leveraging clinical settings, clinical validation, understanding general patterns and decision-making, bias detection, addressing ethical issues, language translations, question answering, information extraction, and business process automation, among others [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. The fine-tuning and retraining capabilities of LLMs empower them to be adapted to specific tasks or behaviours in a predefined manner. Fine-tuning involves taking an already trained language model and retraining its existing weights and bias values using a new dataset specific to a particular domain. This process allows the LLM to be customised and refined for tasks beyond its original training, enhancing its applicability in specific contexts [64]. This process results in a new model that is more tailored and focused on the specified domain. In existing literature, it is frequently observed that LLMs are fine-tuned with a particular emphasis on engineering and research-related fields. This targeted fine-tuning ensures that the model is adept at handling tasks and generating content specifically relevant to the intricacies of these domains [46].
@@ 2.5 Large Action Models
Despite the primarily text-based operations common in LLMs, challenges persist in their ability to manipulate tasks and operations effectively. To address this limitation, recent efforts have introduced innovative concepts such as the Large Action Model (LAM) by the Rabbit research team [^1] https://www.rabbit.tech/research. In this approach, the base model is trained to comprehend sequences of human-provided actions and commands, enabling it to execute these actions accordingly. Similarly, Microsoft introduced the concept of Visualisation-of-Thought (VoT) aimed at integrating human cognitive abilities, specifically the creation of mental images, into the model [65]. Through this approach, it has been demonstrated that MLLMs excel in visual tasks, thereby enabling the extension of action capabilities within a LAM to any LLM. These advancements signify promising directions toward enhancing the practical applicability and versatility of language models across various domains.
@@ 3 Capabilities of Large Language Models
This section focuses on the abilities and capabilities of Language Model Models (LLMs) as outlined in Section 2.4. Additionally, it discusses the currently available fine-tuned LLMs that exhibit potential for application in DF.
3.1 Programming/Coding
The ability to generate source code within a specific context is a crucial proficiency inherent to a language model [66]. Xu et al. [67] conducted a systematic evaluation of six LLMs for code generation across twelve diverse programming languages. The benchmarking process involved the use of the HumanEval benchmark and an evaluation dataset designed to assess the functional correctness of programs generated by an LLM [68]. Initially, the dataset was used to benchmark Code, a GPT language model fine-tuned with publicly available code sourced from GitHub [68]. The Mostly Basic Programming Problems (MBPP) stands as another benchmark, comprising 974 programming tasks. It serves as a frequently used evaluation dataset for LLMs specialising in code-related tasks [69]. Several LLMs explicitly trained for code generation include Code LLaMA, CodeGen, StarCoder, PanGu-Coder, PanGu-Coder2, WizardCoder, InCoder 6B, CodeGen-Mono 16B, Code-Davinci-001, Code-Davinci-002, PaLM-Coder-540B, CodeT5+, InstructCodeT5+, GPT-4 with Reflexion, CodeGeeX, AlphaCode, Santa-Coder [70, 71, 72, 73, 74, 75, 33, 76, 77, 78, 79, 80, 81, 82, 83, 4, 84]. A higher value for both HumanEval and MBPP indicates greater precision in code generation for a given task. For detailed information, refer to Table 1, which presents the counts for HumanEval and MBPP, along with the trained parameter size for each LLM. Even with the HumanEval benchmarking rates, Liu et al. [85] raised questions about the accuracy of the functionality of the generated code. To address this concern, they proposed EvalPlus, an automated test generation engine, to verify the functional correctness of the LLM-generated code.* Estimated parameter count as value is not officially released [86].
Table 1: Trained parameter count, Humaneval and MBPP count for specifically trained LLMs for code generation.
@@ 3.1 Programming/Coding
CodeGen 16.1B 29.28 35.28
CodeGen-MONO 16.1B 29.7 42.4
StarCoder 15.5B 40.8 49.5
PanGu-Coder 2.6B 27.78 23
@@
InstructCodeT5+ 16B 35.0 -
GPT-4 with Reflexion 1.76T* 91.0 77.1
@@ 3.2 Vision Assistance
Traditional vision assistant systems face limitations in image processing or recognition, as they are typically trained on fixed types of datasets. However, with the emergence of LLMs, this paradigm has shifted towards utilising raw text as the source of supervision [87, 88]. Research on visual recognition language models is experiencing exponential growth, with the number of models surpassing 1,500 in 2023 [89]. Radford et al. [87] introduced a novel method called Contrastive Language-Image Pre-training (CLIP). This method is efficient and capable of performing a wide range of tasks during pre-training. It enables the training of a model to learn a shared representation space for both images and text, facilitating a deeper understanding of the relationships between the two modalities. Ramesh et al. [90] proposes a model for text-to-image generation, capable of generating images as combinations derived from textual input or sentences. Moreover, with the model named Generating Images with Large Language Models (GILL), it becomes feasible to generate text, retrieve images, generate novel images, and interleave the results into coherent multimodal dialogues [91]. VisionLLM is a framework leveraging LLMs for diverse vision tasks with unified language instruction, demonstrating generality and flexibility [92]. It incorporates a language-guided image tokeniser and an LLM-based task decoder, capable of handling open-ended tasks based on provided language instructions [92].
@@
Visual instruction tuning leverages language-only models, such as GPT-4, to generate multimodal language-image instruction following data. This data is then utilised to instruction-tune large multimodal models, such as Large Language and Vision Assistant (LLaVA) [93, 94, 7]. The open-source LLaVA project introduces an end-to-end trained model, integrating a vision encoder with an LLM. Notably, LLaVA showcases multimodal chat capabilities. LLaVa exhibits the capability to engage with images, providing detailed descriptions and responding to queries with a reported accuracy of 92.53% [93]. This demonstrates its effectiveness in understanding and generating contextually relevant information about visual content [93]. MiniGPT-4 is an open-source, powerful visual instruction-tuned LLM, and it demonstrates versatility by generating stories and poems inspired by provided images and teaching users how to cook based on visual cues from food photos. This showcases its ability to understand and respond creatively to diverse visual stimuli [95].
@@
Position-enhanced Visual Instruction Tuning (PVIT) represents an extended version of Multimodal Large Language Models (MLLMs). It facilitates region-level encoding in an image, enabling the model to discern and identify information within specific regions [96]. Like PVIT, GPT-4 Region of Interest (ROI) is an instruction-tuned LLM capable of extracting information such as colour, shape, material, and action from a specified region [97]. This model enables users to interact with both language and drawing bounding boxes to indicate the area of interest within an image [97]. Other MLLMs, such as Visual ChatGPT, InternGPT, Flamingo, BLIP-2, and Kosmos, are noted in the literature for their capacity to assist users in visual-related information [98, 99, 100, 101, 102, 103, 104].
@@
Video information is gaining prominence in vision assistance, and Zhao et al. [105] has introduced a novel approach to automatically narrate lengthy videos using LLMs. UniVL is another language pre-trained model designed for both multimodal understanding and generation. It is capable of retrieving a video segment based on text descriptions, generating captions for given video clips, segmenting a video according to a provided text input, and conducting multimodal sentiment analysis of a video segment [106]. VidIL and MOV are additional MLLMs with similar capabilities, demonstrating proficiency in video classification and video-language operations such as video captioning, video question answering, video caption retrieval, and video future event prediction [107, 108].
@@
These MLLMs adhere to a shared task set, encompassing visual question answering, visual captioning, visual common-sense reasoning, visual generation, multimodal affective computing, visual retrieval, vision-language navigation, multimodal machine translation, visual question generation, and visual dialoguing, as summarised in Table 2 [109, 110].
@@ 3.3 Conversation
Specific LLMs are trained explicitly for meaningful and coherent dialogues with humans. An example is Dialogue Generative Pre-trained Transformer (DialoGPT), a fine-tuned model trained on 174 million conversations from Reddit [111]. DialoGPT exhibits the ability to provide human-like answers in tested conversations [111]. Dettmers et al. [112] introduced a fine-tuning mechanism for LLMs named Quantised Pre-trained Language Model into Low-Rank Adapters (QLoRA). This allows the fine-tuning of large-parameter LLMs with low training costs. They introduced Guanaco, a fine-tuned LLM with 65 billion parameters, which achieved a performance level of 99.3%. Falcon-180B and Falcon-40B represent another set of open-source LLMs with 180 billion and 40 billion parameters. These models are trained to communicate in multiple languages, allowing users to engage in conversations in languages other than English [113]. To evaluate the accuracy of human-like dialogue systems, Ou et al. [114] proposed a dialogue evaluation benchmark named DialogBench, comprising 12 dialogue tasks to assess the capabilities of LLMs. In their evaluation, they assessed 28 pre-trained and instruction-tuned LLMs, demonstrating that GPT-4, ChatGPT, and KwaiYii-13B-Chat emerged as the top three models for conversations in domains related to daily life and professional knowledge.
@@ 3.4 Prompt Engineering
Achieving quality outputs from LLMs often relies on providing well-crafted, meaningful, and precise input queries, known as input prompts. However, even human-defined natural language instructions may not consistently yield the best results. Prompt engineering is a methodology that involves carefully defining and instructing LLMs to generate more accurate and desirable outputs. Through thoughtful refinement of input prompts, prompt engineering aims to enhance the performance and effectiveness of LLMs in generating outputs that align more closely with user expectations and requirements [56, 115]. This plays a crucial role in biasing LLMs towards specific domains or topics, enabling a more targeted and nuanced response. By carefully crafting prompts, users can guide LLMs to delve deeper into the nuances of their queries, leading to more accurate and relevant outputs. This approach enhances the model’s responsiveness to specific areas of interest, allowing users to fine-tune and tailor their interactions with the LLM for more precise and meaningful outcomes. Prompt engineering with LLMs is employed across various sectors, including but not limited to medical, engineering, construction, and healthcare [116, 117].
ChainForge is an open-source Graphical User Interface (GUI) tool developed specifically for prompt engineering and hypothesis testing derived from LLMs that can be used in the above-mentioned fields to generate accurate and swift outputs [118].
@@ 3.5 Autonomous Agents
The evolution of LLMs, with their capacity to generate information and communicate in a manner resembling human interaction, has led to the development of autonomous agents. The expectation is that these agents will effectively execute a wide array of tasks, capitalising on the human-like capabilities inherent in LLMs [119]. These autonomous agents follow a four-stage architecture, encompassing profiling, memory, planning, and action. Profiling defines the agent’s role, privileges, domain, and expertise [119]. Memory stores information about tasks and profile data relevant to the environment. Planning involves breaking down given tasks into subtasks and solving them individually. The action stage is the final phase where all decisions and subtasks are translated into actions executed by the agent. Zhang et al. [120] developed a framework designed to facilitate collaboration between AGI agents and humans. This framework enables planning and communication for specific tasks, leveraging the capabilities of LLMs Similarly, AgentSims, ToolBench, GameGPT, ChatDev, Voyager, and RecMind represent a diverse array of autonomous AGI agents developed with distinct goals and objectives [121, 122, 123, 124, 125, 126]. Certainly, AutoGen stands out as a multiagent framework with the capability to autonomously perform tasks or collaborate with human feedback. This flexibility makes it a versatile tool for various applications [127].
In addition to the AutoGen framework, AgentLite, Camel and CrewAI are each similar LLM-based agent framework architectures. These platforms are distinguished by their support for task decomposition, multi-agent orchestration, and adaptable reasoning. Notably, AgentLite and CrewAI facilitate work delegation functionalities, augmenting their utility in various operational contexts [128, 129].

### v3

@@ 2.4 Specifically Trained LLMs
The transformer model and the self-attention mechanism have paved the way for researchers to train language models on trillions of tokens with billions of parameters. Several LLMs have been trained and harnessed, each designed with specific capabilities for diverse fields such as security [33], chemistry [50, 51], engineering [52], medicine [53], business [54], tourism [55], and language-related applications [53]. These models are used in tasks ranging from detecting security threats [33], analysing data and generating synthetic actions to teaching [56], code generation [52, 57], structured query generation [54, 58], planning [54], assisting in medical education [53], clinical decision-making [59], leveraging clinical settings [60], clinical validation [61], understanding general patterns and decision-making [54], bias detection [56], addressing ethical issues [62], language translations [63], question answering [56], information extraction [64], and business process automation [54], among others [53]. The fine-tuning and retraining capabilities of LLMs enable them to be adapted to specific tasks or behaviours in a predefined manner. Fine-tuning involves taking an already trained language model and retraining its existing weights and bias values using a new dataset specific to a particular domain. This process allows the LLM to be customised and refined for tasks beyond its original training, enhancing its applicability in specific contexts [65]. This process results in a new model that is more tailored and focused on the specified domain. In the existing literature, it is frequently observed that LLMs are fine-tuned with a particular emphasis on engineering and research-related fields. This targeted fine-tuning ensures that the model is adept at handling tasks and generating content specifically relevant to the intricacies of these domains [53].
2.5 Multi-modal Large Language Models (MLLMs)
Unlike traditional LLMs, which are trained on text data, Multi-modal Large Language Models (MLLMs) are designed to process and interpret image-based data alongside text. These models can extract and analyse information within images or videos, integrating visual and textual data to enhance comprehension and analysis [66]. The application of MLLMs is rapidly expanding across fields such as digital forensics, where they can be used to analyse images of documents or identify visual anomalies. Section 3.2 provides a discussion on the potential and diverse applications of MLLMs in various domains.
@@ 2.6 Large Action Models
While LLMs excel in text generation and processing, they struggle with complex task manipulation and operational control, especially when moving from language understanding to action execution. This limitation arises because their core design emphasises prediction and generation over direct task execution. To overcome this shortcoming, recent research has introduced innovative approaches, such as the Large Action Model (LAM) developed by the Rabbit research team. LAM extends the capabilities of LLMs by incorporating action-based operations [^3] https://www.rabbit.tech/research. These LAMs can mimic human routines such as scheduling meetings with given instructions, sending emails, ordering taxis, and handling complex tasks such as making reservations for a whole trip. In this approach, the base model is trained to comprehend sequences of human-provided actions and commands, allowing it to execute these actions and tasks accordingly. Similarly, Microsoft introduced the concept of Visualisation-of-Thought (VoT) aimed at integrating human cognitive abilities, specifically the creation of mental images, into the model [67]. Through this approach, it has been demonstrated that MLLMs excel in visual tasks, thereby enabling the extension of action capabilities within an LAM to any LLM. These advancements signify promising directions toward enhancing the practical applicability and versatility of language models across various domains.
@@ 3 Capabilities of Large Language Models
This section focusses on the abilities and capabilities of Language Model Models (LLMs) as outlined in Section 2.4. This section also discusses the currently available fine-tuned LLMs that exhibit potential for application in DF. Although considered too broad for this article, Zhao et al. [68] provide a detailed generic overview of LLMs, their operation, and how they are trained and fine-tuned.
3.1 Programming/Code Generation
The ability to generate source code within a specific context is a crucial skill inherent in a language model [69]. Xu et al. [70] conducted a systematic evaluation of six LLMs for code generation in 12 different programming languages. The benchmarking process employed the HumanEval benchmark, along with a tailored evaluation dataset designed to assess the functional correctness of the programs generated by an LLM [71]. The benchmark comprises a set of coding problems in which the model is tasked with generating Python functions. Each problem is accompanied by a prompt and a set of unit tests that verify whether the generated code produces the expected output. This facilitates a systematic evaluation to generate both syntactically correct and functionally accurate programs. Using this dataset, it is possible to measure performance on real-world coding tasks, as well as its ability to generate solutions that satisfy functional requirements.
The Mostly Basic Programming Problems (MBPP) is another benchmark comprising 974 programming tasks. It serves as a frequently used evaluation dataset for LLMs specialising in code-related tasks [72]. Several LLMs explicitly trained for code generation include Code LLaMA [73], CodeGen [74], StarCoder [75], PanGu-Coder [76], PanGu-Coder2 [76], WizardCoder [77], InCoder 6B [78], CodeGen-Mono 16B [74], Code-Davinci-001 [79], Code-Davinci-002 [79], PaLM-Coder-540B [37], CodeT5+ [80, 81], InstructCodeT5+ [82, 81], GPT-4 with Reflexion [83], CodeGeeX [84], AlphaCode [85], Santa-Coder [86], CodeFuse-13B [87], Codex [88], WaveCoder [89]. A higher value for both HumanEval and MBPP indicates greater precision in code generation for a given task. For detailed information, refer to Table 1, which presents the counts for HumanEval and MBPP, along with the trained parameter size for each LLM. A higher score for both HumanEval and MBPP indicates greater precision in code generation for a given task.
Table 1 presents the scores for HumanEval and MBPP for each of the code generation LLMs mentioned above, along with the trained parameter size for each LLM. Four generic LLMs, or Mixture-of-Experts (MoE) models, are also included in Table 1: o1-mini [90], GPT-4 with Reflexion, DeepSeek-V3 and DeepSeek-V3-Base [8]. These are included as these are the top 4 best performing models for HumanEval despite them being MoE models.* Estimated parameter count as value is not officially released [91].
Table 1: Trained parameter count, HumanEval and MBPP scores for LLM based code generation (ordered by HumanEval score).
@@ 3.1 Programming/Code Generation
o1-mini 100B 97.6 93.9
GPT-4 with Reflexion 1.76T* 91.0 77.1
DeepSeek-V3 671B 85.6 -
DeepSeek-V3-Base 671B 65.2 75.4
@@
CodeGen-MONO 16.1B 29.7 42.4
CodeGen 16.1B 29.28 35.28
Codex-12B 12B 28.81 -
PanGu-Coder 2.6B 27.78 23
@@ 3.2 Vision Assistance
Traditional vision assistant systems face limitations in image processing or recognition, as they are typically trained on fixed types of datasets. However, with the emergence of LLMs, this paradigm has changed to the use of raw text as a source of supervision [92, 93]. Research on visual recognition language models is experiencing exponential growth, with the number of models exceeding 1,500 in 2023 [94]. Radford et al. [92] introduced a novel method called Contrastive Language Image Pre-training (CLIP). This method is efficient and capable of performing a wide range of tasks during pre-training. It enables a model to learn a shared representation space for both images and text, facilitating a deeper understanding of the relationships between the two modalities. Ramesh et al. [95] proposes a model for text-to-image generation, capable of generating images as combinations derived from textual input or sentences. Moreover, with the model named Generating Images with Large Language Models (GILL), it becomes feasible to generate text, retrieve images, generate novel images, and interleave the results into coherent multimodal dialogues [96]. VisionLLM is a framework leveraging LLMs for diverse vision tasks with unified language instruction, demonstrating generality and flexibility [97]. It incorporates a language-guided image tokeniser and an LLM-based task decoder, capable of handling open-ended tasks based on provided language instructions [97].
@@
Visual instruction tuning leverages language-only models, such as GPT-4, to generate multimodal language-image instruction following data. This data is then utilised to instruction-tune large multimodal models, such as Large Language and Vision Assistant (LLaVA) [98, 99, 6]. The open source LLaVA project introduces an end-to-end trained model, integrating a vision encoder with an LLM. Notably, LLaVA showcases multimodal chat capabilities. LLaVa has the capability to interact with images, provide detailed descriptions and respond to queries with a reported accuracy of 92.53% [98]. This shows its effectiveness in understanding and generating contextually relevant information on visual content [98]. MiniGPT-4 is an open-source, powerful visual instruction-tuned LLM, and it demonstrates versatility by generating stories and poems inspired by provided images and teaching users how to cook based on visual cues from food photos. This showcases its ability to understand and respond creatively to various visual stimuli [100].
@@
Position-Enhanced Visual Instruction Tuning (PVIT) represents an extended version of Multimodal Large Language Models (MLLMs). It facilitates region-level encoding in an image, enabling the model to discern and identify information within specific regions [101]. This model enables users to interact with both the language and drawing the bounding boxes to indicate the area of interest within an image [101]. Other MLLMs, such as Visual ChatGPT [92], InternGPT [102], Flamingo [103], BLIP-2 [104], and Kosmos [105], are noted in the literature for their ability to assist users in visual-related information.
@@
Video information is gaining prominence in vision assistance, and Zhao et al. [106] has introduced a novel approach to automatically narrate lengthy videos using LLMs. UniViLM is another language pre-trained model designed for both multimodal understanding and generation. It is capable of retrieving a video segment based on text descriptions, generating captions for given video clips, segmenting a video according to a provided text input, and performing multimodal sentiment analysis of a video segment [107]. VidIL and LLaViLo are additional MLLMs with similar capabilities, demonstrating proficiency in video classification and video-language operations such as video captioning, video question answering, video caption retrieval, and prediction of future video events [108, 109].
@@
These MLLMs adhere to a shared task set, that includes visual question answering, visual captioning, visual common-sense reasoning, visual generation, multimodal affective computing, visual retrieval, vision language navigation, multimodal machine translation, visual question generation and visual dialoguing, as summarised in Table 2 [110, 111].
@@ 3.3 Conversation
Specific LLMs are trained explicitly for meaningful and coherent dialogues with humans. An example is Dialogue Generative Pre-trained Transformer (DialoGPT), a fine-tuned model trained on 174 million Reddit conversations [112]. DialoGPT exhibits the ability to provide human-like answers in tested conversations [112]. Dettmers et al. [113] introduced a fine-tuning mechanism for LLMs named Quantised Pre-trained Language Model into Low-Rank Adapters (QLoRA). This allows for the fine-tuning of large-parameter LLMs with low training costs. They introduced Guanaco, a fine-tuned LLM with 65 billion parameters, which achieved a performance level of 99.3%. Falcon-180B and Falcon-40B represent another set of open-source LLMs with 180 billion and 40 billion parameters. These models are trained to communicate in multiple languages, allowing users to engage in conversations in languages other than English [114]. To evaluate the accuracy of human-like dialogue systems, Ou et al. [115] proposed a dialogue evaluation benchmark named DialogBench, which consists of 12 dialogue tasks to assess the capabilities of LLMs. In their evaluation, they assessed 28 pre-trained and instruction-tuned LLMs, demonstrating that GPT-4, ChatGPT, and KwaiYii-13B-Chat emerged as the top three models for conversations in domains related to daily life and professional knowledge.
@@ 3.4 Prompt Engineering
Achieving quality outputs from LLMs often relies on providing well-crafted, meaningful, and precise input queries, known as input prompts. However, even human-defined natural language instructions may not consistently yield the best results. Prompt engineering is a methodology that involves carefully defining and instructing LLMs to generate more accurate and desirable outputs. Through thoughtful refinement of input prompts, prompt engineering aims to enhance the performance and effectiveness of LLMs in generating output that align more closely with user expectations and requirements [116, 117]. This plays a crucial role in biasing LLMs toward specific domains or topics, enabling a more targeted and nuanced response. By carefully crafting prompts, users can guide LLMs to dive deeper into the nuances of their queries, leading to more accurate and relevant outputs. This approach enhances the model’s responsiveness to specific areas of interest, allowing users to fine-tune and tailor their interactions with the LLM for more precise and meaningful outcomes. Prompt engineering with LLMs is employed in various sectors, including, but not limited to, medical, engineering, construction, and healthcare [118, 119].
ChainForge is an open source Graphical User Interface (GUI) tool developed specifically for prompt engineering and hypothesis testing derived from LLMs that can be used in the aforementioned fields to generate accurate and quick output [120].
Although prompt engineering is a necessity in generating the desired output from an LLM, the results can still be biassed based on the wording and phrasing provided by the user. Since the effectiveness of prompts depends on the user’s proficiency in English, the output may vary significantly depending on the exact requirements of the user and how the LLM interprets these prompts. In addition, the prompts can unintentionally reinforce existing biases within the model’s training data, potentially skewing the results. Therefore, prompt engineering must be approached carefully and methodically to minimise misinterpretation and maximise output relevance.
@@ 3.5 Autonomous Agents
The evolution of LLMs, with their ability to generate information and communicate in a manner that resembles human interaction, has led to the development of autonomous agents. The expectation is that these agents will effectively perform a wide range of tasks, taking advantage of the human-like capabilities inherent in LLMs [121]. These autonomous agents follow a four-stage architecture that includes profiling, memory, planning, and action. Profiling defines the agent’s role, privileges, domain, and expertise [121]. Memory stores information on tasks and profile data relevant to the environment. Planning involves breaking down given tasks into subtasks and solving them individually. The action stage is the final phase where all decisions and subtasks are translated into actions executed by the agent. Zhang et al. [122] developed a framework designed to facilitate collaboration between GenAI agents and humans. This framework enables planning and communication for specific tasks, leveraging the capabilities of LLMs Similarly, AgentSims [123], ToolBench [124], GameGPT [125], ChatDev [126], Voyager [127], and RecMind [128] represent a diverse array of autonomous GenAI agents developed with distinct goals and objectives. Certainly, AutoGen stands out as a multiagent framework with the capability to autonomously perform tasks or collaborate with human feedback. This flexibility makes it a versatile tool for various applications [129].
In addition to the AutoGen framework, AgentLite [130], Camel [131] and CrewAI [132] are each similar LLM-based agent framework architectures. These platforms are distinguished by their support for task decomposition, multi-agent orchestration, and adaptable reasoning. In particular, AgentLite and CrewAI facilitate work delegation functionalities, increasing their utility in various operational contexts.